OpenStudy (suwhitney):

Log/Expo. Question... 8^x=1/2^(-3/x) Please help! :)

OpenStudy (suwhitney):

\[8^{x}=\frac{ 1 }{ 2^{^{-\frac{3 }{ x }}} }\]

OpenStudy (anonymous):

Right side is the same as \(\large 2^{3/x}\).

OpenStudy (anonymous):

is it like this one? 8^x=(1/2)^(-3/x)

OpenStudy (anonymous):

And the left side is the same as \(\large 2^{3x}\) because 8=2^3. Can you handle it from there, @suwhitney ?

OpenStudy (suwhitney):

Sorry, my computer is being slow.. im looking over everything

OpenStudy (suwhitney):

So I understand that that is what the right side is but then do I make it \[8=(2^{3})^{^{\frac{ 3 }{ x }}}\]??

OpenStudy (anonymous):

use the left hand side as is as 8^x

OpenStudy (suwhitney):

I dont change it to 2^3?

OpenStudy (anonymous):

8^x =8^1/x then whats next

OpenStudy (suwhitney):

x=1/x ?

OpenStudy (suwhitney):

well wouldnt it be x=3/x ?

OpenStudy (anonymous):

x=1/x is correct.

OpenStudy (anonymous):

Hint: This is a quadratic equation so will have two solutions.

OpenStudy (suwhitney):

1 and -1?

OpenStudy (anonymous):

Those are the only two that work. It might be worthwhile to check both of those in the original equation to verify.

OpenStudy (anonymous):

yeheeyyy ... :D you got it

OpenStudy (suwhitney):

Ok, thanks so much both of yall :)

OpenStudy (anonymous):

Recap: \[\large 8^x=\frac{1}{2^{-3/x}}\] \[\large \rightarrow 2^{3x}=2^{3/x}\] \[\large \rightarrow {3x}={3/x}\] \[\large \rightarrow {x}={1/x}\] \[\large \rightarrow x^2=1, \space x= \pm1\]

OpenStudy (suwhitney):

Thank you so much!! :)

OpenStudy (anonymous):

You're welcome. Don't let these nasty-looking equations intimidate you. Break them down one step at a time!

OpenStudy (suwhitney):

I'll try not to :)